OpenStudy (anonymous):

A jogger runs into the countryside at a rate of 10mph. He returns along the same route at 6mph. If the total trip is 1 hour 36 minutes how far did he jog?

OpenStudy (anonymous):

@tcarroll010 ?

OpenStudy (anonymous):

i think its 6m

OpenStudy (anonymous):

I have the answer as being 12 miles but don't know how to figure it out

OpenStudy (anonymous):

both the distances are the same so by taking the distance x, total time is given by x/10+x/6 = 1.6 hours thats how i did it

OpenStudy (anonymous):

Rt=d so how would I set this up?

OpenStudy (anonymous):

Time=dist/speed

OpenStudy (anonymous):

Time there+time back=1.6

OpenStudy (anonymous):

I set it up as 96=d/16 Then divided 96 by 16 and got 6 then multiplied by 2 since 6 would have been half

OpenStudy (anonymous):

d/10+d/6=1.6

OpenStudy (anonymous):

12 is suppose to be answer I got twelve by doing it